# REP: Resource-Efficient Prompting for Rehearsal-Free Continual Learning

### v4

@@ 4.3 Additional studies
We further assess the efficacy of REP by applying it to CL methods not covered by Table˜1, including two additional non-prompting rehearsal-free methods: such as SLCA [40] and RanPAC [24]. Additionally, we compare REP-L2P with two adapter-based methods, including Online-LoRA [39] and ADAM [15]. More discussions are provided in Appendix˜H and Appendix˜J, respectively.
@@ Acknowledgements
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants (No. RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH); No. RS-2025-02304554, Efficient and Scalable Framework for AI Heterogeneous Cluster Systems; No. RS-2022-II220290, Visual Intelligence for Space-Time Understanding and Generation), the National Research Foundation of Korea (NRF) grants (No. 2021R1A2C2012195; No. RS-2024-00354947), and the Electronics and Telecommunications Research Institute (ETRI) grant (No. 25ZS1100, Research on High-Performance Computing to Overcome Limitations of AI), all funded by the Korean government (MSIT).
@@ B.1 Prompt-based methods for REP
L2P positions prompts at the first layer of the transformer architecture [38]. These prompts are learnable parameters that dynamically evolve with the training process. The mechanism begins with a prompt pool $P=\{p_{1},p_{2},...,p_{m}\}\subset\mathbb{R}^{L_{p}\times D}$.
@@
DualPrompt leverages prompts at multiple layers of the transformer architecture [37]. It introduces a general prompt $g\in\mathbb{R}^{L_{g}\times D}$ and a set of task-specific prompts $E=\{e_{1},e_{2},...,e_{T}\}\subset\mathbb{R}^{L_{e}\times D}$. These prompts are incorporated at specified layers in the transformer model.
@@
CODA-Prompt decomposes learnable prompts into components and uses an attention mechanism from a pre-trained ViT model to select relevant prompts [31]. Instead of a single prompt, CODA-Prompt learns a set of prompt components $P=\{P_{1},P_{2},...,P_{M}\}$. The final prompt $p$ is calculated as a weighted sum:
@@
where $P^{n}$, $K^{n}$, and $A^{n}$ are new components, and $\lambda$ balances the orthogonality loss [31].
HiDe-Prompt decomposes the CL objective into three parts—within-task prediction (WTP), task-identity inference (TII), and task-adaptive prediction (TAP) [33]. It manages uninstructed/instructed representations by modeling each class $c$ with approximate distributions (e.g., Gaussian means $\boldsymbol{\mu}_{c}$). The overall combined loss is composed with WTP loss ($\mathcal{L}_{\mathrm{WTP}}$), TII loss ($\mathcal{L}_{\mathrm{TII}}$), TAP loss ($\mathcal{L}_{\mathrm{TAP}}$) and a contrastive regularization loss ($\mathcal{L}_{\mathrm{CR}}$):
@@
ConvPrompt generates layer-wise prompts via convolution on a shared embedding matrix $\mathrm{SE}$ [30]. For a layer $\ell$ and head $h$, it produces $M$ prompt components $\mathrm{PC}^{K}_{\ell,h,m}$, $\mathrm{PC}^{V}_{\ell,h,m}$ by convolving $\mathrm{SE}^{K}_{\ell,h}$ and $\mathrm{SE}^{V}_{\ell,h}$ with $m$th prompt generator $G_{\ell,h,m}$. The final prompt is computed by weighting these components via a projection network and prompt keys:
@@
OVOR employs a single prompt throughout all tasks, in contrast to methods that maintain a large prompt pool [13]. To further refine decision boundaries, it introduces virtual outlier regularization (VOR), which generates synthetic outliers $\mathbf{v}\in\mathcal{D}_{\mathrm{Outlier}}$ to penalize overconfident predictions outside the training distribution. OVOR’s loss is described as:
@@
LAE reframes prompts as an instance of parameter-efficient fine-tuning (PEFT), incorporating additional modules (Adapters or LoRA) to mitigate forgetting [7]. It accumulates multi-task knowledge through online/offline PET modules and ensembles them at inference. Concretely, it defines an online PET $\boldsymbol{\theta}_{\mathrm{pet}}^{\text{on}}$ for new tasks and an offline PET $\boldsymbol{\theta}_{\mathrm{pet}}^{\text{off}}$ for older tasks:
@@ B.2 Hyperparameters and configurations
For all prompt-based CL methods, we employ the ADAM optimizer [15] with hyperparameters $\beta_{1}=0.990$ and $\beta_{2}=0.999$, following their original implementation. For prompt selection, we set the prompt pool size to 10 and the prompt length to 5, following the original implementations of L2P, DualPrompt, and HiDe-Prompt. For CODA-Prompt, we implement a cosine-decay learning rate strategy described in [31], while other prompt-based baselines use a fixed learning rate of $1.875\times 10^{-3}$ by default.
We consistently use the mini-batch size of 16 for all methods to maintain a uniform computational load for each training iteration. For fair comparisons, each method performs 1875, 1080, and 2583 iterations per task for Split CIFAR-100 [18], Split ImageNet-R [12], and Split PlantDisease [25], respectively. This setup ensures that among prompt-based methods, iteration time itself correlates linearly with energy usage, as training wall-clock time similarly reflects energy consumption, as discussed in the main paper.
@@ Appendix D Preliminary empirical studies
In this studies, we uncover the cost-accuracy trade-offs to highlight the impact of each baseline method on the accuracy gained relative to its training cost. Here, we limit the memory to a maximum of 8GB. We employ L2P [38], DualPrompt [37], and CODA-Prompt [31] as representative ViT-based prompting methods. To demonstrate the effectiveness of prompting methods, we also incorporate two state-of-the-art CNN-based methods, such as BudgetCL [27] and MEMO [43]. Both BudgetCL and MEMO improve model accuracy by leveraging spare memory to store and replay past samples, with MEMO additionally storing model checkpoints from history. These methods also use ImageNet pre-trained models as backbones: ResNet-10 (RN10; 5M), ResNet-18(RN18; 11M), ResNet-26 (RN26; 14M), ResNet-34 (RN34;22M), and ResNet-101 (RN101; 43M) models.
For BudgetCL and MEMO, we adopt the SGD optimizer with a learning rate scheduling and a weight decay, following their original papers [27, 43]. Both methods construct mini-batches of 16 samples, with 8 samples from the new task and 8 from prior tasks in the replay buffer. We adopt the setup in [43] and train for 200 epochs across all datasets. Thus, BudgetCL and MEMO perform more iterations per epoch with a larger replay buffer, resulting in longer training time and higher energy usage. For fair comparisons, we group methods by memory usage.
@@
Memory efficiency.  CNN-based methods are often considered more memory-efficient because they use backbones a magnitude smaller than ViT models. Thus, it is commonly believed that ample memory can be allocated to replay samples or the past model checkpoints [43], which help improve model performance. However, in CL, which involves typical training iterations, a substantial amount of memory must be allocated for feature maps [21], leaving little room for memory buffers. For instance, for a device memory range of 1–4GB in Table˜11, the backbone of MEMO/RN18 consumes only 43MB of memory (6.5% of ViT-L), but feature maps occupy as much as 1.7GB (more than ViT-based methods). This causes MEMO/RN18 to be short on memory for the replay buffer, dramatically degrading model accuracy, as observed in Fig.˜5.
@@ Appendix E Extended ablation studies
Next, we perform algorithm validation on Split CIFAR-100 (10 tasks), for which we present the results in Table˜13. Consistent with the Split ImageNet-R results, using static token merging (ToMe) [2] or progressive layer dropping (PLD) [41] instead of adaptive token merging (AToM) or adaptive layer dropping (ALD) negatively affects model accuracy. This experiment further corroborates the efficacy of our proposed algorithms for optimizing resource usage without compromising the model accuracy in CL setup. We further extend our analysis from Table˜4 by evaluating two additional prompt-based CL methods, including HiDe-Prompt (denoted as REP-HiDeP) and ConvPrompt (denoted as REP-ConvP), with results detailed in Table˜14. These results show a trend consistent with Table˜4.
@@ Appendix G Additional datasets
We also expand our experiments on REP by incorporating an additional dataset: Split CUB-200 [32], which is designed for fine-grained image classification. Split CUB-200 contains 5,994 bird images categorized into 200 classes, which are split into 5 tasks, each having 40 classes. As shown in Table˜17, integrating REP achieves resource efficiency by bringing significant reductions in both training time and memory usage.
@@ Appendix H Supplementary evaluations
Self-supervised pre-training.  We evaluate REP with HiDe-Prompt [33] under two self-supervised pre-training paradigms, including iBOT21K [44] and DINO [3], using the ViT-B backbone on Split ImageNet-R (10 tasks). As shown in Table˜18, REP reduces training time and memory usage by 14–18%, with only 0.6–0.8% of accuracy drop while forgetting remains improved.
@@
Comparison with analytic learning-based CL method.  We evaluate F-OAL [45] as the representative analytic learning-based CL method on Split CIFAR-100 (10 tasks). Table˜19 demonstrates that REP-L2P can achieve compatible accuracy, with shorter training time. Since F-OAL is an online method, a perfect 1:1 comparison is tricky. Nonetheless, these results suggest that REP can be competitive with F-OAL’s energy efficiency while maintaining similar accuracy. Each method has distinct advantages and disadvantages depending on the context.
@@ Appendix I Extended main results including forgetting metric
Guided by the hyperparameter tuning strategies used in DualPrompt [37] and CODA-Prompt [31], we fine-tuned key hyperparameters for our ALD approach using REP-L2P. Our focus is primarily on optimizing the threshold parameter ($\alpha$) and the adjustment factor ($\tau$), which are crucial for ALD’s adaptability and efficiency. We conducted a hyperparameter search using 5-fold cross-validation, for which we designate 20% of our training dataset as a validation set. Table˜21 summarizes the results.
We further evaluate the impact of varying prompt-selection-related hyperparameters using REP-L2P on subsets of 5-datasets dataset introduced in [38]. Table˜22 shows that the primary factor driving computation costs is prompt length, as it introduces more learnable parameters (e.g., with a pool size of 10, increasing prompt length from 5 to 10 increases memory usage by 500 MB and training time by 14%). In contrast, increasing the prompt pool size has almost no impact on training time and memory usage.
@@
Miro [22] introduces a dynamic approach for fine-tuning key design parameters of rehearsal-based CL methods to achieve high model accuracy while simultaneously reducing energy costs, i.e., high cost-effectiveness. Miro’s methodology centers on identifying the suitable memory size within the device’s capacity to accommodate both old and new samples for training. We compare REP-L2P with Miro directly on a reference edge device, NVIDIA Jetson TX2. This device is equipped with a 256-core NVIDIA Pascal GPU, 8GB of RAM, and a 32GB eMMC 5.1 drive. The RAM is a single unified memory shared by the CPU and GPU. We report energy usage as joules (J) obtained by multiplying power by time. Power usage is measured by reading the built-in sensor values provided by Jetson devices for the GPU, RAM, CPU, and I/O connection.
@@
In this experiment, we explore energy-accuracy trade-offs of REP-L2P and Miro. Figure˜7 visually represents the comparison results. In each graph, the x-axis signifies total energy usage (lower is better), while the y-axis signifies final average accuracy (higher is better). Hence, a more cost-effective strategy is positioned closer to the upper-left corner of the graph. Regarding Miro, we maintain the hyperparameters as suggested in the original work [22] but incorporate the use of pre-trained ResNet-18 (RN18; 11M), ResNet-34 (RN34; 22M), and ResNet-50 (RN50; 25M)[11] models instead of non-pre-trained ones to enhance overall performance. For REP-L2P, we vary the number of training iterations per task insertion to match the energy usage of Miro variants with different ResNet models.

### v6

@@ 4.3 Additional studies
We further assess the efficacy of REP by applying it to CL methods not covered by Table˜1, including two additional non-prompting rehearsal-free methods: such as SLCA [slca] and RanPAC [ranpac]. Additionally, we compare REP-L2P with two adapter-based methods, including Online-LoRA [onlinelora] and ADAM [adam]. More discussions are provided in Appendix˜H and Appendix˜J, respectively.
@@ Acknowledgements
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants (No. RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH); No. RS-2024-00459797, Development of ML Compiler Framework for On-device AI; No. RS-2025-02304554, Efficient and Scalable Framework for AI Heterogeneous Cluster Systems; No. RS-2022-II220290, Visual Intelligence for Space-Time Understanding and Generation), the National Research Foundation of Korea (NRF) grants (No. RS-2024-00337559; No. RS-2024-00354947), and the Electronics and Telecommunications Research Institute (ETRI) grant (No. 25ZS1100, Research on High-Performance Computing to Overcome Limitations of AI), all funded by the Korean government (MSIT).
@@ B.1 Prompt-based methods for REP
L2P positions prompts at the first layer of the transformer architecture [l2p]. These prompts are learnable parameters that dynamically evolve with the training process. The mechanism begins with a prompt pool $P=\{p_{1},p_{2},...,p_{m}\}\subset\mathbb{R}^{L_{p}\times D}$.
@@
DualPrompt leverages prompts at multiple layers of the transformer architecture [dualp]. It introduces a general prompt $g\in\mathbb{R}^{L_{g}\times D}$ and a set of task-specific prompts $E=\{e_{1},e_{2},...,e_{T}\}\subset\mathbb{R}^{L_{e}\times D}$. These prompts are incorporated at specified layers in the transformer model.
@@
CODA-Prompt decomposes learnable prompts into components and uses an attention mechanism from a pre-trained ViT model to select relevant prompts [codap]. Instead of a single prompt, CODA-Prompt learns a set of prompt components $P=\{P_{1},P_{2},...,P_{M}\}$. The final prompt $p$ is calculated as a weighted sum:
@@
where $P^{n}$, $K^{n}$, and $A^{n}$ are new components, and $\lambda$ balances the orthogonality loss [codap].
HiDe-Prompt decomposes the CL objective into three parts—within-task prediction (WTP), task-identity inference (TII), and task-adaptive prediction (TAP) [hidep]. It manages uninstructed/instructed representations by modeling each class $c$ with approximate distributions (e.g., Gaussian means $\boldsymbol{\mu}_{c}$). The overall combined loss is composed with WTP loss ($\mathcal{L}_{\mathrm{WTP}}$), TII loss ($\mathcal{L}_{\mathrm{TII}}$), TAP loss ($\mathcal{L}_{\mathrm{TAP}}$) and a contrastive regularization loss ($\mathcal{L}_{\mathrm{CR}}$):
@@
ConvPrompt generates layer-wise prompts via convolution on a shared embedding matrix $\mathrm{SE}$ [convp]. For a layer $\ell$ and head $h$, it produces $M$ prompt components $\mathrm{PC}^{K}_{\ell,h,m}$, $\mathrm{PC}^{V}_{\ell,h,m}$ by convolving $\mathrm{SE}^{K}_{\ell,h}$ and $\mathrm{SE}^{V}_{\ell,h}$ with $m$th prompt generator $G_{\ell,h,m}$. The final prompt is computed by weighting these components via a projection network and prompt keys:
@@
OVOR employs a single prompt throughout all tasks, in contrast to methods that maintain a large prompt pool [ovor]. To further refine decision boundaries, it introduces virtual outlier regularization (VOR), which generates synthetic outliers $\mathbf{v}\in\mathcal{D}_{\mathrm{Outlier}}$ to penalize overconfident predictions outside the training distribution. OVOR’s loss is described as:
@@
LAE reframes prompts as an instance of parameter-efficient fine-tuning (PEFT), incorporating additional modules (Adapters or LoRA) to mitigate forgetting [lae]. It accumulates multi-task knowledge through online/offline PET modules and ensembles them at inference. Concretely, it defines an online PET $\boldsymbol{\theta}_{\mathrm{pet}}^{\text{on}}$ for new tasks and an offline PET $\boldsymbol{\theta}_{\mathrm{pet}}^{\text{off}}$ for older tasks:
@@ B.2 Hyperparameters and configurations
For all prompt-based CL methods, we employ the ADAM optimizer [adam] with hyperparameters $\beta_{1}=0.990$ and $\beta_{2}=0.999$, following their original implementation. For prompt selection, we set the prompt pool size to 10 and the prompt length to 5, following the original implementations of L2P, DualPrompt, and HiDe-Prompt. For CODA-Prompt, we implement a cosine-decay learning rate strategy described in [codap], while other prompt-based baselines use a fixed learning rate of $1.875\times 10^{-3}$ by default.
We consistently use the mini-batch size of 16 for all methods to maintain a uniform computational load for each training iteration. For fair comparisons, each method performs 1875, 1080, and 2583 iterations per task for Split CIFAR-100 [cifar], Split ImageNet-R [imagenet_r], and Split PlantDisease [plantdisease], respectively. This setup ensures that among prompt-based methods, iteration time itself correlates linearly with energy usage, as training wall-clock time similarly reflects energy consumption, as discussed in the main paper.
@@ Appendix D Preliminary empirical studies
In this studies, we uncover the cost-accuracy trade-offs to highlight the impact of each baseline method on the accuracy gained relative to its training cost. Here, we limit the memory to a maximum of 8GB. We employ L2P [l2p], DualPrompt [dualp], and CODA-Prompt [codap] as representative ViT-based prompting methods. To demonstrate the effectiveness of prompting methods, we also incorporate two state-of-the-art CNN-based methods, such as BudgetCL [budgetcl] and MEMO [memo]. Both BudgetCL and MEMO improve model accuracy by leveraging spare memory to store and replay past samples, with MEMO additionally storing model checkpoints from history. These methods also use ImageNet pre-trained models as backbones: ResNet-10 (RN10; 5M), ResNet-18(RN18; 11M), ResNet-26 (RN26; 14M), ResNet-34 (RN34;22M), and ResNet-101 (RN101; 43M) models.
For BudgetCL and MEMO, we adopt the SGD optimizer with a learning rate scheduling and a weight decay, following their original papers [budgetcl, memo]. Both methods construct mini-batches of 16 samples, with 8 samples from the new task and 8 from prior tasks in the replay buffer. We adopt the setup in [memo] and train for 200 epochs across all datasets. Thus, BudgetCL and MEMO perform more iterations per epoch with a larger replay buffer, resulting in longer training time and higher energy usage. For fair comparisons, we group methods by memory usage.
@@
Memory efficiency.  CNN-based methods are often considered more memory-efficient because they use backbones a magnitude smaller than ViT models. Thus, it is commonly believed that ample memory can be allocated to replay samples or the past model checkpoints [memo], which help improve model performance. However, in CL, which involves typical training iterations, a substantial amount of memory must be allocated for feature maps [zico], leaving little room for memory buffers. For instance, for a device memory range of 1–4GB in Table˜11, the backbone of MEMO/RN18 consumes only 43MB of memory (6.5% of ViT-L), but feature maps occupy as much as 1.7GB (more than ViT-based methods). This causes MEMO/RN18 to be short on memory for the replay buffer, dramatically degrading model accuracy, as observed in Fig.˜5.
@@ Appendix E Extended ablation studies
Next, we perform algorithm validation on Split CIFAR-100 (10 tasks), for which we present the results in Table˜13. Consistent with the Split ImageNet-R results, using static token merging (ToMe) [token_merging] or progressive layer dropping (PLD) [progressive_drop] instead of adaptive token merging (AToM) or adaptive layer dropping (ALD) negatively affects model accuracy. This experiment further corroborates the efficacy of our proposed algorithms for optimizing resource usage without compromising the model accuracy in CL setup. We further extend our analysis from Table˜4 by evaluating two additional prompt-based CL methods, including HiDe-Prompt (denoted as REP-HiDeP) and ConvPrompt (denoted as REP-ConvP), with results detailed in Table˜14. These results show a trend consistent with Table˜4.
@@ Appendix G Additional datasets
We also expand our experiments on REP by incorporating an additional dataset: Split CUB-200 [CUB200], which is designed for fine-grained image classification. Split CUB-200 contains 5,994 bird images categorized into 200 classes, which are split into 5 tasks, each having 40 classes. As shown in Table˜17, integrating REP achieves resource efficiency by bringing significant reductions in both training time and memory usage.
@@ Appendix H Supplementary evaluations
Self-supervised pre-training.  We evaluate REP with HiDe-Prompt [hidep] under two self-supervised pre-training paradigms, including iBOT21K [ibot] and DINO [dino], using the ViT-B backbone on Split ImageNet-R (10 tasks). As shown in Table˜18, REP reduces training time and memory usage by 14–18%, with only 0.6–0.8% of accuracy drop while forgetting remains improved.
@@
Comparison with analytic learning-based CL method.  We evaluate F-OAL [foal] as the representative analytic learning-based CL method on Split CIFAR-100 (10 tasks). Table˜19 demonstrates that REP-L2P can achieve compatible accuracy, with shorter training time. Since F-OAL is an online method, a perfect 1:1 comparison is tricky. Nonetheless, these results suggest that REP can be competitive with F-OAL’s energy efficiency while maintaining similar accuracy. Each method has distinct advantages and disadvantages depending on the context.
@@ Appendix I Extended main results including forgetting metric
Guided by the hyperparameter tuning strategies used in DualPrompt [dualp] and CODA-Prompt [codap], we fine-tuned key hyperparameters for our ALD approach using REP-L2P. Our focus is primarily on optimizing the threshold parameter ($\alpha$) and the adjustment factor ($\tau$), which are crucial for ALD’s adaptability and efficiency. We conducted a hyperparameter search using 5-fold cross-validation, for which we designate 20% of our training dataset as a validation set. Table˜21 summarizes the results.
We further evaluate the impact of varying prompt-selection-related hyperparameters using REP-L2P on subsets of 5-datasets dataset introduced in [l2p]. Table˜22 shows that the primary factor driving computation costs is prompt length, as it introduces more learnable parameters (e.g., with a pool size of 10, increasing prompt length from 5 to 10 increases memory usage by 500 MB and training time by 14%). In contrast, increasing the prompt pool size has almost no impact on training time and memory usage.
@@
Miro [miro] introduces a dynamic approach for fine-tuning key design parameters of rehearsal-based CL methods to achieve high model accuracy while simultaneously reducing energy costs, i.e., high cost-effectiveness. Miro’s methodology centers on identifying the suitable memory size within the device’s capacity to accommodate both old and new samples for training. We compare REP-L2P with Miro directly on a reference edge device, NVIDIA Jetson TX2. This device is equipped with a 256-core NVIDIA Pascal GPU, 8GB of RAM, and a 32GB eMMC 5.1 drive. The RAM is a single unified memory shared by the CPU and GPU. We report energy usage as joules (J) obtained by multiplying power by time. Power usage is measured by reading the built-in sensor values provided by Jetson devices for the GPU, RAM, CPU, and I/O connection.
@@
In this experiment, we explore energy-accuracy trade-offs of REP-L2P and Miro. Figure˜7 visually represents the comparison results. In each graph, the x-axis signifies total energy usage (lower is better), while the y-axis signifies final average accuracy (higher is better). Hence, a more cost-effective strategy is positioned closer to the upper-left corner of the graph. Regarding Miro, we maintain the hyperparameters as suggested in the original work [miro] but incorporate the use of pre-trained ResNet-18 (RN18; 11M), ResNet-34 (RN34; 22M), and ResNet-50 (RN50; 25M)[resnet] models instead of non-pre-trained ones to enhance overall performance. For REP-L2P, we vary the number of training iterations per task insertion to match the energy usage of Miro variants with different ResNet models.